OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? x+3y=12 -4x+3y=-3 A. (0, 4) B. (6, 2) C. (3, 3) D. (0, −1)

OpenStudy (anonymous):

@TheSavior

OpenStudy (anonymous):

C 3+3(3)=12 3+9=12 12=12 -4x+3y=-3 -4(3)+ 3(3)=-3 -12 + 9 = -3 -3=-3

OpenStudy (anonymous):

It's not A

OpenStudy (anonymous):

I know :)

OpenStudy (anonymous):

It's not B

OpenStudy (anonymous):

I know :)

OpenStudy (anonymous):

It's C

OpenStudy (anonymous):

It's not D

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

It was correct:)

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Good Job!!! (if you did figure it out yourself xD)